# OrderChain: A General Prompting Paradigm to Improve the Ordinal Understanding Ability of MLLM

### v1

@@ Abstract
Despite the remarkable progress of multimodal large language models (MLLMs), they continue to face challenges in achieving competitive performance on ordinal regression (OR; a.k.a. ordinal classification). To address this issue, this paper presents OrderChain, a novel and general prompting paradigm that improves the ordinal understanding ability of MLLMs by specificity and commonality modeling. Specifically, our OrderChain consists of a set of task-aware prompts to facilitate the specificity modeling of diverse OR tasks and a new range optimization Chain-of-Thought (RO-CoT), which learns a commonality way of thinking about OR tasks by uniformly decomposing them into multiple small-range optimization subtasks. Further, we propose a category recursive division (CRD) method to generate instruction candidate category prompts to support RO-CoT automatic optimization. Comprehensive experiments show that a Large Language and Vision Assistant (LLaVA) model with our OrderChain improves baseline LLaVA significantly on diverse OR datasets, e.g., from 47.5% to 93.2% accuracy on the Adience dataset for age estimation, and from 30.0% to 85.7% accuracy on the Diabetic Retinopathy dataset. Notably, LLaVA with our OrderChain also remarkably outperforms state-of-the-art methods by 27% on accuracy and 0.24 on MAE on the Adience dataset. To our best knowledge, our OrderChain is the first work that augments MLLMs for OR tasks, and the effectiveness is witnessed across a spectrum of OR datasets.
@@ 1 Introduction
Large language models (LLMs), e.g., GPT3 [6], LLaMA [46], Gemini [45], and Qwen [2], have shown unprecedented capabilities in understanding human languages and solving practical problems such as scientific question answering and code generation. When integrated with visual encoders, large language models can be upgraded to multimodal large language models (MLLMs), like LLaVA [30] and GPT-4V [37], which can achieve the ability similar to human visual intelligence and tackle visual understanding tasks [52, 50]. Despite these advances, the potential of MLLMs on order understanding for ordinal regression tasks is not yet well explored.
@@
Figure 3: The overall framework of our proposed OrderChain for improving the ordinal understanding of MLLMs (e.g., LLaVA). Given a query image, the MLLM is instructed to: (I) Format the raw input of the task as preliminary identification, (II) generate a definition of the task as domain knowledge prompts, (III) divide target labels or previous coarse prediction into candidate subsets as prompts for more refined prediction, and (IV) after obtaining the final prediction, return the response.
@@
To tackle these two challenges, this work explores how the ordinal understanding capability of MLLMs can be realized by developing a new Chain-of-Thought method, called OrderChain. OrderChain is motivated by three crucial considerations. First, domain-related knowledge needs to be utilized to guide the MLLM to extract critical imaging features and make predictions more effectively. Second, for different OR tasks, the knowledge on the number and range of candidate category labels needs to be injected to mitigate the out-of-bound predictions made by the MLLM. Both these two types of knowledge can help the MLLM to model the specificity of different OR tasks. Third, inspired by the idea of decomposition, OR tasks can be decomposed into sub-interval classification tasks. Through continuous decomposition and classification, the candidate range can be gradually optimized to obtain the final prediction. This insight is applicable to all tasks whose categories have an order, which can help the MLLM to model the commonality of OR tasks. Based on these considerations, our proposed OrderChain consists of the following key components:
(1) A comprehensive set of task-aware prompts, aiming to guide the MLLM for specificity modeling. The task-aware prompts include domain knowledge prompts and category feature prompts; the former offers some prior domain knowledge to facilitate the MLLM’s understanding of the target task, and the latter provides the number and range of candidate category label information to urge the MLLM to make more reasonable predictions.
(2) A range optimization Chain-of-Thought (RO-CoT) that enables the MLLM to solve OR tasks in a progressive manner. As illustrated in Fig. 3, our RO-CoT is designed to (1) format the raw input with a suitable instruction template, (2) generate a definition of the task as domain knowledge prompts, (3) divide target labels or previous coarse prediction into multiple smaller candidate subsets as category feature prompts with a category recursive division method, and use these prompts with domain knowledge prompts for more refined prediction, and (4) return the final response until the final prediction is made. Given an image and an OR task query, our RO-CoT can automatically decompose it into multimodal range optimization subtasks and refine the predicted candidate subsets progressively.
(3) A category recursive division method that computes and divides the labels of target tasks into coarse-to-fine candidate categories. This method can automatically divide the previous prediction of RO-CoT into multiple refined candidate subsets, which are processed as part of category feature prompts to support the continuous range optimization of RO-CoT.
@@
Compared to traditional deep learning (DL) methods, Our OrderChain allows MLLMs to train a unified model for all OR tasks. Extensive experiments on OR datasets of various domains show the effectiveness of our OrderChain and its components. Notably, OrderChain achieves state-of-the-art performance in facial age estimation tasks, remarkably improving accuracy to 93.2% (a 27% improvement). In other domains of OR tasks, LLaVA with our OrderChain yields improvement of $\sim$12to $\sim$56compared to baseline LLaVA, showing highly competitive performance.
@@ 2.2 Chain-of-Thought (CoT)
Figure 4: An example of the Category Recursive Division (CRD) process on the facial age estimation task. Assume that the total number of categories is 80 and CRD uses a trinomial balanced tree, that is, each node has three child nodes (subsets) and the difference in the number of categories per child node (sub candidate sets) is minimal. CRD first divides the entire range (80) of categories into three subsets with new category numbers 27, 27 and 26. It repeats this process until the number of categories for each subset is less than or equal to 3. Upon the last division, the final prediction is obtained. $N$ at each node denotes the number of categories within the node (candidate set).
@@ 3.2 Range Optimization Chain-of-Thought
    Formatting the Query to a Template: Transforming individualized raw query inputs to a specific template of common OR tasks.
@@
    Loop Iteration in Each Range Optimization Step:
@@
      Generating category feature prompts by the category recursive division (CRD) method (discuss in Sec. 3.3): Given the initial optional categories, CRD will continue to divide these categories into multiple consecutive subsets with a smaller range as new candidate categories for the next step.
@@
      Proposing a thought based on the task-aware prompts (discuss in Sec. 3.4): The task-aware prompts, consisting of domain knowledge prompts and category feature prompts, guide the MLLM to propose a thought – which new candidate subset that the category of the image belongs to.
@@ 3.3 Category Recursive Division Method
Figure 5: An example of OrderChain for facial age estimation on the Adience dataset. Brown: Domain knowledge prompts. Green: Description prompts. Blue: Instruction prompts.
@@ 4.1 Datasets and Setup
   Method Accuracy (%) $\uparrow$ MAE $\downarrow$
   Poisson [5] 77.1 0.38
   MT [41] 82.8 0.36
   SORD [11] 78.2 0.73
   POE [26] 80.5 0.30
   CIG [8] 83.3 0.30
   Ord2Seq [47] 84.2 0.25
   LLaVA-1.5 (zero-shot) [30] 3.9 12.1
   LLaVA-1.5 (baseline) [30] 30.0 0.99
   LLaVA-1.5 + OrderChain 85.7 0.23
Table 4: Accuracy and MAE comparison on the DR dataset.
@@ 4.2 Experimental Results
Facial Age Estimation. Fig. 5 gives an example of OrderChain for facial age estimation on the Adience dataset, and Table 1 shows comparison results with state-of-the-art (SOTA) methods on the Adience dataset. One can see that the zero-shot version (training-free) and baseline version (LoRA fine-tuning) of LLaVA-1.5 perform worse than known supervised SOTA methods. Specifically, the zero-shot version of LLaVA-1.5 attains only 17.6% accuracy and baseline version of LLaVA -1.5 perform better but merely obtains a mediocre performance. In contrast, LLaVA with our proposed OrderChain achieves 93.2% accuracy and 0.12 MAE, outperforming SOTA methods L2RCLIP [48] by a remarkable margin of $\sim$27improvement in accuracy and 0.24 reduction in MAE. This indicates that when integrated with our OrderChain, the MLLM is instructed to think of the facial age estimation problem step by step with a few smaller refined subtasks, which allows the MLLM to learn the internal ordinal relationships between categories and effectively estimate the ages of faces with a wide range and a large number of categories. Our Ordchain achieves milestone performance on age estimation tasks, which further demonstrates the effectiveness of OrderChain in improving the ordinal understanding of MLLM.
Image Aesthetics Assessment. Table 3 shows the results on the Image Aesthetics dataset. We find that the MLLM with our OrderChain does not achieve optimal performance, which we believe is due to the highly subjective nature of the labels made by human raters. The subjective differences between different people, and even between people and MLLMs for the definition of beauty, may be significant. Especially for the relatively low performance on the ‘People’ category, we suspect it is due to the pre-training imposed on the MLLM that tends to praise rather than demean people. In other relatively more objective categories, LLaVA with our OrderChain achieves higher performance, and thus the overall performance is mainly affected by the ‘People’ category. On the other hand, our proposed OrderChain remarkably improves vanilla LLaVA to a competitive level, demonstrating the effectiveness of our OrderChain.
Diabetic Retinopathy Grading. Table 4 shows the results on the DR dataset. Note that the DR dataset is unbalanced since the sample number decreases sharply as the severity DR level increases. We observe that the known methods yield poor performances, possibly due to the unbalanced data. Especially, SORD [11], which is a modality-specific method utilizing modified soft labels, suffers serious errors in MAE. Worse, the zero-shot and baseline versions of LLaVA attain horrible performance. In comparison, LLaVA with our proposed OrderChain still maintains competitive performances, achieving an Accuracy of 85.7% and an MAE of 0.23, which greatly outperforms the baselines and the other order learning methods, showing that our approach has better robustness on unbalanced data. We believe that this is due to the category range division process of RO-CoT, which also enables better positive-negative distinction. That is, unlike one positive class against the other negative classes in previous work, it turns to (e.g.) classifying the first two categories against the last three categories in the first CoT step of OrderChain for the DR dataset (5 categories in total). In this way, the classification in a step is more category-balanced, which helps to better deal with unbalanced data. Moreover, our OrderChain provides a promising way to grade medical diseases of the OR type by a unified MLLM.
@@ 4.3 Ablation Study
5 Limitations and Future Work
Our approach still has some limitations. The first limitation is that on OR tasks whose labels are highly subjective, such as aesthetic assessment, the performance of MLLM still has room for improvement. In the future, we will further explore stronger domain or rater knowledge for MLLM to understand label subjectivity. Another limitation is that although our model possesses a general ordinal understanding, the zero-shot performance on datasets of unseen domains may suffer. Future work will focus on exploring few-shot methods that need merely a few samples for the adaptation of our unified MLLM model on a new domain task.
@@ 6 Conclusions
In this paper, we presented a novel and general prompting paradigm, OrderChain, to improve the ordinal understanding of MLLMs for ordinal regression. We first pointed out two major reasons for the limited performance of vanilla MLLMs, i.e., lack of specificity modeling and commonality modeling. We adopted a range optimization Chain-of-Thought to learn a commonality way of thinking about ordinal regression tasks and task-aware prompts to inject task-specific information into MLLMs. We also introduce a category recursive division tool to generate refined candidate category subsets for supporting range optimization. Extensive experiments showed that our OrderChain significantly improves the performance of the MLLM and achieves optimal performance in most ordinal regression tasks, especially on the facial age estimation task, with 27% overall accuracy improvement and 0.24 MAE reduction, demonstrating that OrderChain can effectively improve the ordinal understanding of MLLMs and provides a promising paradigm to build a unified ordinal regression MLLM.

### v3

# OrderChain: Towards General Instruct-Tuning for Stimulating the Ordinal Understanding Ability of MLLM

@@ Abstract
Despite the remarkable progress of multimodal large language models (MLLMs), they continue to face challenges in achieving competitive performance on ordinal regression (OR; a.k.a. ordinal classification). To address this issue, this paper presents OrderChain, a novel and general prompting paradigm that improves the ordinal understanding ability of MLLMs by specificity and commonality modeling. Specifically, our OrderChain consists of a set of task-aware prompts to facilitate the specificity modeling of diverse OR tasks and a new range optimization Chain-of-Thought (RO-CoT), which learns a commonality way of thinking about OR tasks by uniformly decomposing them into multiple small-range optimization subtasks. Further, we propose a category recursive division (CRD) method to generate instruction candidate category prompts to support RO-CoT automatic optimization. Comprehensive experiments show that LLaVA model with our OrderChain improves baseline LLaVA significantly on diverse OR datasets, e.g., from 47.5% to 93.2% accuracy on the Adience dataset for age estimation, and from 30.0% to 85.7% accuracy on the Diabetic Retinopathy dataset. Notably, LLaVA with our OrderChain also remarkably outperforms state-of-the-art methods by 27% on accuracy and 0.24 on MAE on the Adience dataset. To our best knowledge, our OrderChain is the first work that augments MLLMs for OR tasks, and the effectiveness is witnessed across a spectrum of OR datasets. Project Page: https://order-chain.github.io/.
@@ 1 Introduction
With the success of Large language models (LLMs), e.g., GPT3 [6], LLaMA [49], Gemini [48], and Qwen [3], researchers have been enhancing these models by incorporating visual understanding capabilities and led to emergence of multimodal large language models (MLLMs), such as LLaVA [34] and GPT-4V [40] and Qwen-VL [3, 4], exhibiting capabilities comparable to human intelligence in visual understanding tasks [55, 53]. Despite these advances, the potential of MLLMs on order understanding for ordinal regression tasks is not yet well explored.
@@
Figure 3: The overall framework of our proposed OrderChain for improving the ordinal understanding of Multi-modal LLMs (e.g., LLaVA).
@@
To tackle these two challenges, this work investigates how the ordinal understanding capability of MLLMs can be realized by developing a novel Chain-of-Thought method, called OrderChain. OrderChain is driven by three pivotal motivations. First, domain-related knowledge needs to be utilized to guide the MLLM to extract critical imaging features and make predictions more effectively. Second, for different OR tasks, the knowledge on the number and range of candidate category labels needs to be injected to mitigate the out-of-bound predictions made by the MLLM. Both these two types of knowledge can help the MLLM to model the specificity of different OR tasks. Third, inspired by the idea of decomposition, OR tasks can be decomposed into sub-interval classification tasks. Through continuous decomposition and classification, the candidate range can be gradually optimized to obtain the final prediction. This insight is applicable to all tasks whose categories have an order, which can help the MLLM to model the commonality of OR tasks. Based on these considerations, our proposed OrderChain consists of the following key components: (1) Customized task-aware prompts to guide MLLMs in specificity modeling. These include domain knowledge prompts, which provide prior knowledge to help understand the task, and category feature prompts, which offer information on the number and range of candidate labels to enable more accurate predictions. (2) A Range Optimization Chain-of-Thought (RO-CoT) for progressive ordinal regression. As shown in Fig. 3, RO-CoT transforms inputs into instruction templates, generates task definitions as domain knowledge prompts, recursively divides labels or coarse predictions into smaller subsets as category feature prompts, and iteratively refines predictions until a final response is achieved. This approach decomposes visual rating tasks into subtasks for gradual refinement. (3) A category recursive division method for coarse-to-fine candidate label generation. This method automatically splits RO-CoT predictions into refined subsets, which are used as category feature prompts to support continuous optimization in the prediction process.
@@
Compared to traditional deep learning (DL) methods, Our OrderChain allows MLLMs to train a unified model for all OR tasks. Extensive experiments on OR datasets of various domains show the effectiveness of our OrderChain and its components. Notably, OrderChain achieves state-of-the-art performance in facial age estimation tasks, remarkably improving accuracy to 93.2% (a 27% improvement). In other domains of OR tasks, LLaVA with our OrderChain yields improvement of $\sim$12% to $\sim$56% compared to baseline LLaVA, showing highly competitive performance.
@@ 3.2 Range Optimization Chain-of-Thought
    Formatting and Setting Template: Format individualized raw query inputs to a specific template of common OR tasks and set corresponding recursive CoT logic.
@@
    Loop Iteration in Each RO-CoT Step:
@@
      Generating category feature prompts by the category recursive division (CRD) method (discuss in Sec. 3.3): Given the initial optional categories or last intermediate thought, CRD will continue to divide these categories into multiple consecutive subsets with a smaller range as new candidate categories for the next step. These candidate subsets will serve as category feature prompts for instruction.
@@
      Using combined task-aware prompts to instruct MLLM (discuss in Sec. 3.4): The task-aware prompts, combined with default domain knowledge prompts and newly generated category feature prompts, are used to instruct the MLLM to propose a thought which a new candidate subset that the category of the image belongs to.
@@
Figure 4: An example of the Category Recursive Division (CRD) process on the facial age estimation task. Assume that the total number of categories is 80 and CRD uses a trinomial balanced tree, that is, each node has three child nodes (subsets) and the difference in the number of categories per child node (sub candidate sets) is minimal. CRD first divides the entire range (80) of categories into three subsets with new category numbers 27, 27 and 26. It repeats this process until the number of categories for each subset is less than or equal to 3. Upon the last division, the final prediction is obtained. NN at each node denotes the number of categories within the node (candidate set).
@@ 3.3 Category Recursive Division Method
Figure 5: An example of OrderChain for facial age estimation on the Adience dataset. Brown: Domain knowledge prompts. Blue: Description prompts. Green: Instruction prompts.
@@ 4 Experiments
4.1 Datasets and Setup
   Method Accuracy (%) $\uparrow$ MAE $\downarrow$
   Poisson [5] 77.1 0.38
   MT [44] 82.8 0.36
   SORD [11] 78.2 0.73
   POE [29] 80.5 0.30
   CIG [8] 83.3 0.30
   Ord2Seq [50] 84.2 0.25
   LLaVA-1.5 (zero-shot) [34] 3.9 12.1
   LLaVA-1.5 (baseline) [34] 30.0 0.99
   LLaVA-1.5 + OrderChain 85.7 0.23
Table 3: Accuracy and MAE comparison on the DR dataset.
@@ 4.2 Experimental Results
Facial Age Estimation. Fig. 5 gives an example of OrderChain for facial age estimation on the Adience dataset, and Table 1 shows comparison results with state-of-the-art (SOTA) methods on the Adience dataset. One can see that the zero-shot version (training-free) and baseline version (LoRA fine-tuning) of LLaVA-1.5 perform worse than known supervised SOTA methods. Specifically, the zero-shot version of LLaVA-1.5 attains only 17.6% accuracy and baseline version of LLaVA -1.5 perform better but merely obtains a mediocre performance. In contrast, LLaVA with our proposed OrderChain achieves 93.2% accuracy and 0.12 MAE, outperforming SOTA methods L2RCLIP [51] by a remarkable margin of $\sim$27% improvement in accuracy and 0.24 reduction in MAE. This indicates that when integrated with our OrderChain, the MLLM is instructed to think of the facial age estimation problem step by step with a few smaller refined subtasks, which allows the MLLM to learn the internal ordinal relationships between categories and effectively estimate the ages of faces with a wide range and a large number of categories. Our Ordchain achieves milestone performance on age estimation tasks, which further demonstrates the effectiveness of OrderChain in improving the ordinal understanding of MLLM.
Diabetic Retinopathy Grading. Table 3 shows the results on the DR dataset. Note that this dataset is unbalanced since the sample number decreases sharply as the severity DR level increases. Known methods yield poor performances due to the unbalanced data, like SORD [11], which is a modality-specific method utilizing modified soft labels, suffers serious errors in MAE. Worse, the zero-shot and baseline versions of LLaVA attain horrible performance. In comparison, LLaVA with our OrderChain still maintains competitive performances, achieving an Accuracy of 85.7% and an MAE of 0.23, which greatly outperforms the baselines and the other order learning methods, showing that our approach has better robustness on unbalanced data. It is mainly due to the category range division process of RO-CoT, which also enables better positive-negative distinction. Unlike one positive class against the other negative classes in previous work, it turns to classifying the first two categories against the last three categories in the first CoT step of OrderChain for the DR dataset (5 categories in total). In this way, the classification in a step is more category-balanced, which helps to better deal with unbalanced data. Moreover, our OrderChain provides a promising way to grade medical diseases of the OR type by a unified MLLM.
Image Aesthetics Assessment. Table 4 shows the results on the Image Aesthetics dataset. We find that the MLLM with our OrderChain does not achieve optimal performance, which we believe is due to the highly subjective nature of the labels made by human raters. The subjective differences between different people, and even between people and MLLMs for the definition of beauty, may be significant. Especially for the relatively low performance on the ‘People’ category, we suspect it is due to the pre-training imposed on the MLLM that tends to praise rather than demean people. To further illustrate this point, we provide quantitative visualization results as shown in the F sig. 6. Our model successfully predicted the aesthetic scores for the first and third images. However, for the second image, the model makes an incorrect prediction with a higher score, which also reflects the issue we described in the paper: MLLMS have positive biases towards people.. In other relatively more objective categories, LLaVA with our OrderChain achieves higher performance, and thus the overall performance is mainly affected by the ‘People’ category. On the other hand, our proposed OrderChain remarkably improves vanilla LLaVA to a competitive level, demonstrating the effectiveness of our OrderChain.
Figure 6: Quantitative visualization results of some examples of people aesthetics assessment.
@@ 4.3 Ablation Study
4.4 Zero-shot Generalization
To further validate that our OrderChain learned the general way of thinking for ordinal understanding, we conduct inference-only experiments to show the zero-shot performance of baseline LLaVA and our OrderChain, as shown in Tab. 6. We choose three additional datasets including two in-domain datasets (APTOS [21] dataset for DR grading and AVA dataset for aesthetic assessment [18]) and one out-of-domain dataset (KonIQ [19] for image quality assessment). As we can see, our OrderChain can bring a remarkable zero-shot improvement to the baseline LLaVA on both in-domain and out-of-domain datasets, substantiating the generalization ability of OrderChain. Specifically, the significant improvement on the APTOS dataset can be attributed to to the model’s prior training on similar tasks, enabling it to better adapt to the understanding of this specific task. However, even for unseen scenarios (i.e. image quality assessment), our OrderChain can still obtain a 15% improvement compared to baseline LLaVA. This demonstrates that our approach indeed stimulates the ordinal understanding ability of MLLMs, equipping them with a generalized way of thinking for ordinal regression tasks.
   Method Metric APTOS [21] AVA [18] KonIQ [19]
   Baseline ACC $\uparrow$ 23.2 49.6 36.3
MAE $\downarrow$ 1.02 0.56 0.81
   OrderChain ACC $\uparrow$ 70.0 57.7 51.5
MAE $\downarrow$ 0.30 0.44 0.49
Table 6: Zero-shot Performance of baseline LLaVA and our Orderchain on in-domain and out-of-domain datasets.
@@ 5 Conclusions
In this paper, we presented a novel and general prompting paradigm, OrderChain, to improve the ordinal understanding of MLLMs for ordinal regression. We first pointed out two major reasons for the limited performance of vanilla MLLMs, i.e., lack of specificity modeling and commonality modeling. We adopted a range optimization Chain-of-Thought to learn a commonality way of thinking about ordinal regression tasks and task-aware prompts to inject task-specific information into MLLMs. We also introduce a category recursive division tool to generate refined candidate category subsets for supporting range optimization. Extensive experiments showed that our OrderChain significantly improves the performance of the MLLM and achieves optimal performance in most ordinal regression tasks and considerable zero-shot generalization performance, demonstrating that OrderChain can effectively improve the ordinal understanding of MLLMs and provides a promising paradigm to build a unified ordinal regression MLLM.
Acknowledgements
This research was partially supported by National Natural Science Foundation of China under grants No. 82202984, No. 12326612, Zhejiang Key R&D Program of China under grant No. 2024SSYS0026, and the Transvascular lmplantation Devices Research Institute (TIDRI).